Can someone please help me out with this equation please?
A dr office mixed a 75% disinfectant solution with a 25% disinfectant solution. How many liters of each were used to make 20L of a 40% disinfectant solution. Thanks

Answer by ptaylor(2198)


Let x=number of liters of 75% solution
Then 20-x=number of liters of 25% soluton
Now we know that the amount of pure disinfectant in the 75% solution (0.75x) plus the amount of pure disinfectant in the 25% solution(0.25(20-x)) must equal the amount of pure disinfectant in the final mixture (0.40(20)). So our equation to solve is:
0.75x+0.25(20-x)=0.40*20 get rid of parens
0.75x+5-0.25x=8 subtract 5 from both sides
0.75x+5-5-0.25x=8-5 collect like terms
0.50x=3 divide both sides by 0.50
x=6 liters -------------------------amount of 75% solution needed
20-x=20-6=14 liters---------------------amount of 25% solution needed
CK
6*0.75+14*0.25=20*0.40
4.5+3.5=8
8=8

Hope this helps---------------------ptaylor